OpenStudy (anonymous):

WILL AWARD MEDALS The first term of an arithmetic sequence is -5, and the tenth term is 13. Find the common difference. A. 8/9 B. 1.8 c. 2 I think its B, but I am not 100% sure

OpenStudy (anonymous):

\[a,a+d.a+2d,a+3d,...,a+9d\] \[-5,-5+d,-5+2d,-5+3d,...,-5+9d=13\] solve \[-5+9d=13\] for \(d\)

OpenStudy (anonymous):

you do not get \(1.8\) though

rishavraj (rishavraj):

\[a_1 = -5~~~~and ~~~~a_{10} = 13\] u must be knowing tht \[a_n = a_1 + (n - 1)d\] so \[a_{10} = a_1 + (10 - 1)d\] \[13 = -5 + 9d\] solve and get "d"

OpenStudy (anonymous):

Thanks :)